# ARTIST: Improving the Generation of Text-rich Images by Disentanglement

### v1

@@ 3 Methodology
In this work, we propose to mitigate the aforementioned challenges by utilizing separate modules to learn text structure and visual appearance. Furthermore, these two modules are trained separately, making it possible to learn text structure with synthetic data constructed by ourselves, which also tackles the problem of limited data. To ensure disentanglement and prevent information leakage, our text module and visual module take different prompts as inputs. However, it can be inefficient and user-hostile if these prompts have to be manually designed by the user themselves at inference. Fortunately, because of the recent success of large-language models (LLMs) [3, 22, 36, 37], we propose to utilize pre-trained LLMs to infer user’s intention, provide accurate prompts for both modules. With the help of LLM, user’s input can be either precise or vague, leading to better interactive experience.
@@ 3.1 LLM-based Prompt Understanding
TextDiffuser [6] proposes to train a transformer model to extract texts which are expected to be shown on generated images. However, there appears to be a limitation associated with this method: their model is only capable of detecting words enclosed by quotation marks. This happens because most keywords from the training dataset are enclosed by quotation marks inside the captions. A model trained on these samples will lack generalization ability or even be overfitting. Moreover, their model even fails to generate the desired image according to the prompt “a poster of Batman with word Batman on it", as it does not understand that the word “Batman" should be presented on the image. On the contrary, we propose to provide pre-trained LLM with an open-domain prompt and rely on its capability to identify the essentials autonomously. Because of its vast open-domain knowledge, LLM is able to understand user intents better and thus can generalize to complicated scenarios. Given vague prompts, LLM can discern and propose which words or text elements to incorporate, resulting in more coherent and aesthetically appealing suggestions.
@@ 3.2 Learning Text Structure
After constructing both datasets, we train the text module in two stages, inside the latent space of a VAE [14] following Rombach et al. [28]. In the first stage, the diffusion model is trained to take a single bounding box and a target word as inputs then generate an black-white image with the word rendered on it. Then, this module will be fine-tuned on our sentence-level dataset as the second stage so that it can take multiple bounding boxes and multiple s as inputs. Our s are denoted as $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, composed of the words to be generated in the image. Our input mask is an image containing the corresponding bounding boxes $\{m_{1},m_{2},\ldots m_{n}\}$ that indicate the position of each word. During training, the original text-only image and input mask image are first encoded into latent space features $\mathbf{z}$ and $\mathbf{m}$. Then we sample a time step tsimilar-$t\sim$ Uniform $\left(0,T\right)$ and a Gaussian noise $\boldsymbol{\epsilon}$ to corrupt the original feature, yielding $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\overline{\alpha}_{T}$ is the coefficient of the diffusion process introduced in [10]. $\mathbf{z}_{t}$ and $\mathbf{m}$ are concatenated in the feature channel as input for the diffusion model, which will be trained with the diffusion loss between the sampled $\boldsymbol{\epsilon}$ and the predicated noise $\boldsymbol{\epsilon}_{\theta}$:
@@ 3.3 Learning Visual Appearance
During the training of visual module, the text module will be frozen, and only the newly-introduced layers and visual module will be fine-tuned. Differently from the text module described in Section 3.2, whose training text only contains the target words to be rendered, the training data for our visual module has to contain visual descriptions of the image so that the model can successfully learn to generate visual contents based on the user’s input.
Let $\overline{\mathcal{P}}$ be the prompt which contains visual descriptions of the image, $\mathcal{P}$ be the s as defined in Section 3.2. We feed $\mathcal{P}$ and the input mask $\mathbf{m}$ as mentioned in Section 3.2 into the pre-trained text module. The intermediate features from the text module denoted as $\{f_{i}(\mathbf{z}_{t},\mathbf{m},\mathcal{P},t)\}_{i=1}^{k}$ will be injected into the visual module, where $k$ stands for the number of intermediate features. The visual module will be trained with diffusion loss:
@@ 4.1 Implementation Details
Our experiments are conducted on 8 Nvidia A100 GPUs, with Hugging Face Diffusers [39]. Both text and visual modules are initialized from the pre-trained Stable Diffusion [29] checkpoint. The text module is first pre-trained on our word-level synthetic dataset for 400,000 steps, then further fine-tuned for 200,000 steps on our sentence-level synthetic dataset. With the frozen text module, our visual module is trained for 250,000 steps on MARIO-10M dataset [6]. AdamW [18] optimizer is used in all training stages, with a learning rate of 1e-5 and a weight decay of 1e-2. The batch size is set to 128.
We first compare our proposed framework with previous methods including Stable Diffusion [29] (denoted as SD), ControlNet [49], DeepFloyd [35] and TextDiffuser [6] in terms of common metrics such as OCR accuracy and FID. We also fine-tune a Stable Diffusion on MARIO-10M dataset for a more comprehensive comparison. Then, we directly compare our method with the latest approaches like AnyText [38] and TextDiffuser 2 [5] through human evaluation, as it encapsulates the capabilities of the above traditional metrics while also assessing subjective aspects crucial for user experience and practical application effectiveness. Two variants of our proposed framework are evaluated, which are denoted as ARTIST-TD and ARTIST, respectively, indicating whether LLM is utilized. Specifically, ARTIST-TD directly uses the pre-trained transformer from TextDiffuser instead of LLM to obtain bounding boxes and keywords based on input prompts. Thus, comparing ARTIST-TD with TextDiffuser can straightforwardly show the effectiveness of our training strategy, as they share exactly the same layout and keyword conditions.
@@
Metrics SD Fine-tuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@ 4.2 Main Results
Metrics SD Fine-tuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@ 4.3 Human Evaluation
In the section, we compare our method with GlyphControl [46], AnyText [38], Controlnet [49], Textdiffuser [6], and Textdiffuser-2 [5] We referred to Textdiffuser’s survey design for our evaluation. We collected 49 cases, which is three times the number of cases in Textdiffuser’s study. In each case, we provided images generated by the aforementioned methods to the participants and asked them to rate them based on text rendering quality and image-text matching. The rating scores ranged from 1 to 4, with 4 indicating the best. Similar to Textdiffuser’s study, We collected 30 survey responses in total. (The prompts for these 49 cases were sampled from various datasets. We included 7 from ARTIST-Eval, 7 from ChineseDrawText, 2 from DrawBenchText, 7 from DrawTextCreative, 12 from LAIONEval4000, 7 from OpenLibraryEval500, and 7 from TMDBEval500, considering their different dataset sizes. We provide some examples from the survey in the appendix.
@@ 4.4 Ablation Studies
Although a character segmentation mask is used in our implementation following  [6], the proposed ARTIST should lead to improvements no matter whether the character segmentation mask is used or not. To verify this claim, we conduct an ablation study to show the robustness of ARTIST. Specifically, we trained an ARTIST model without the character-level segmentation mask, which is to be compared with the fine-tuned Stable Diffusion model. For all the models, we apply 50 sampling steps and classifier-free guidance [11] of $\omega\in\{2.5,5,7.5,10,15\}$ during inference. In general, a larger $\omega$ leads to better FID and CLIP score, while it may lead to a worse OCR evaluation. The comparison is provided in Figure 6, from which we can conclude that our ARTIST leads to improvement across different settings because it leads to better alignment of the image and the text, image fidelity, and OCR accuracy compared to baselines.

### v2

# ARTIST: Improving the Generation of Text-rich Images with Disentangled Diffusion Models

@@ 3 Methodology
In this work, we propose to mitigate the aforementioned challenges by utilizing separate modules to learn text structure and visual appearance. Furthermore, these two modules are trained separately, making it possible to learn text structure with synthetic data constructed by ourselves, which also tackles the problem of limited data. To ensure disentanglement and prevent information leakage, our text module and visual module take different prompts as inputs. However, it can be inefficient and user-hostile if these prompts have to be manually designed by the user themselves at inference. Fortunately, because of the recent success of large-language models (LLMs) [3, 22, 36, 37], we propose to utilize pretrained LLMs to infer user’s intention, provide accurate prompts for both modules. With the help of LLM, user’s input can be either precise or vague, leading to better interactive experience.
@@ 3.1 LLM-based Prompt Understanding
TextDiffuser [6] proposes to train a transformer model to extract texts which are expected to be shown on generated images. However, there appears to be a limitation associated with this method: their model is only capable of detecting words enclosed by quotation marks. This happens because most keywords from the training dataset are enclosed by quotation marks inside the captions. A model trained on these samples will lack generalization ability or even be overfitting. Moreover, their model even fails to generate the desired image according to the prompt “a poster of Batman with word Batman on it", as it does not understand that the word “Batman" should be presented on the image. On the contrary, we propose to provide pretrained LLM with an open-domain prompt and rely on its capability to identify the essentials autonomously. Because of its vast open-domain knowledge, LLM is able to understand user intents better and thus can generalize to complicated scenarios. Given vague prompts, LLM can discern and propose which words or text elements to incorporate, resulting in more coherent and aesthetically appealing suggestions.
@@ 3.2 Learning Text Structure
After constructing both datasets, we train the text module in two stages, inside the latent space of a VAE [14] following Rombach et al. [28]. In the first stage, the diffusion model is trained to take a single bounding box and a target word as inputs then generate an black-white image with the word rendered on it. Then, this module will be finetuned on our sentence-level dataset as the second stage so that it can take multiple bounding boxes and multiple s as inputs. Our s are denoted as $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, composed of the words to be generated in the image. Our input mask is an image containing the corresponding bounding boxes $\{m_{1},m_{2},\ldots m_{n}\}$ that indicate the position of each word. During training, the original text-only image and input mask image are first encoded into latent space features $\mathbf{z}$ and $\mathbf{m}$. Then we sample a time step tsimilar-$t\sim$ Uniform $\left(0,T\right)$ and a Gaussian noise $\boldsymbol{\epsilon}$ to corrupt the original feature, yielding $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\overline{\alpha}_{T}$ is the coefficient of the diffusion process introduced in [10]. $\mathbf{z}_{t}$ and $\mathbf{m}$ are concatenated in the feature channel as input for the diffusion model, which will be trained with the diffusion loss between the sampled $\boldsymbol{\epsilon}$ and the predicated noise $\boldsymbol{\epsilon}_{\theta}$:
@@ 3.3 Learning Visual Appearance
During the training of visual module, the text module will be frozen, and only the newly-introduced layers and visual module will be finetuned. Differently from the text module described in Section 3.2, whose training text only contains the target words to be rendered, the training data for our visual module has to contain visual descriptions of the image so that the model can successfully learn to generate visual contents based on the user’s input.
Let $\overline{\mathcal{P}}$ be the prompt which contains visual descriptions of the image, $\mathcal{P}$ be the s as defined in Section 3.2. We feed $\mathcal{P}$ and the input mask $\mathbf{m}$ as mentioned in Section 3.2 into the pretrained text module. The intermediate features from the text module denoted as $\{f_{i}(\mathbf{z}_{t},\mathbf{m},\mathcal{P},t)\}_{i=1}^{k}$ will be injected into the visual module, where $k$ stands for the number of intermediate features. The visual module will be trained with diffusion loss:
@@ 4.1 Implementation Details
Our experiments are conducted on 8 Nvidia A100 GPUs, with Hugging Face Diffusers [39]. Both text and visual modules are initialized from the pretrained Stable Diffusion [29] checkpoint. The text module is first pretrained on our word-level synthetic dataset for 400,000 steps, then further finetuned for 200,000 steps on our sentence-level synthetic dataset. With the frozen text module, our visual module is trained for 250,000 steps on MARIO-10M dataset [6]. AdamW [18] optimizer is used in all training stages, with a learning rate of 1e-5 and a weight decay of 1e-2. The batch size is set to 128.
We first compare our proposed framework with previous methods including Stable Diffusion [29] (denoted as SD), ControlNet [49], DeepFloyd [35] and TextDiffuser [6] in terms of common metrics such as OCR accuracy and FID. We also finetune a Stable Diffusion on MARIO-10M dataset for a more comprehensive comparison. Then, we directly compare our method with the latest approaches like AnyText [38] and TextDiffuser 2 [5] through human evaluation, as it encapsulates the capabilities of the above traditional metrics while also assessing subjective aspects crucial for user experience and practical application effectiveness. Two variants of our proposed framework are evaluated, which are denoted as ARTIST-TD and ARTIST, respectively, indicating whether LLM is utilized. Specifically, ARTIST-TD directly uses the pretrained transformer from TextDiffuser instead of LLM to obtain bounding boxes and keywords based on input prompts. Thus, comparing ARTIST-TD with TextDiffuser can straightforwardly show the effectiveness of our training strategy, as they share exactly the same layout and keyword conditions.
@@
Metrics SD finetuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@ 4.2 Main Results
Metrics SD finetuned SD ControlNet DeepFloyd TextDiffuser ARTIST-TD ARTIST
@@ 4.3 Human Evaluation
In the section, we compare our method with GlyphControl [46], AnyText [38], Controlnet [49], Textdiffuser [6], and Textdiffuser-2 [5] in terms of human evalution in Table 4. We referred to Textdiffuser’s survey design for our evaluation. We collected 49 cases, which is three times the number of cases in Textdiffuser’s study. In each case, we provided images generated by the aforementioned methods to the participants and asked them to rate them based on text rendering quality and image-text matching. The rating scores ranged from 1 to 4, with 4 indicating the best. Similar to Textdiffuser’s study, We collected 30 survey responses in total. (The prompts for these 49 cases were sampled from various datasets. We included 7 from ARTIST-Eval, 7 from ChineseDrawText, 2 from DrawBenchText, 7 from DrawTextCreative, 12 from LAIONEval4000, 7 from OpenLibraryEval500, and 7 from TMDBEval500, considering their different dataset sizes.) We provide some examples from the survey in the appendix.
@@ 4.4 Ablation Studies
Although a character segmentation mask is used in our implementation following  [6], the proposed ARTIST should lead to improvements no matter whether the character segmentation mask is used or not. To verify this claim, we conduct an ablation study to show the robustness of ARTIST. Specifically, we trained an ARTIST model without the character-level segmentation mask, which is to be compared with the finetuned Stable Diffusion model. For all the models, we apply 50 sampling steps and classifier-free guidance [11] of $\omega\in\{2.5,5,7.5,10,15\}$ during inference. In general, a larger $\omega$ leads to better FID and CLIP score, while it may lead to a worse OCR evaluation. The comparison is provided in Figure 6, from which we can conclude that our ARTIST leads to improvement across different settings because it leads to better alignment of the image and the text, image fidelity, and OCR accuracy compared to baselines.